Use the CRAMMER'S RULE to solve these systems of linear equations.

5x+6y+7z=40

2x+4y+2z=34

x+3y+5z=30


1
Expert's answer
2021-05-07T14:33:43-0400

Given:

"\left\{
\begin{array}{ c c c }
5x+6y+7z & = & 40 \\ 
2x+4y+2z & = & 34\\ 
x+3y+5z & = & 30 \\
\end{array}\right."

Solution:

CRAMMER'S RULE 

"\det\left|
\begin{array}{ c c c }
a_{11} & a_{12} & a_{13} \\ 
a_{21} & a_{22} & a_{23}\\ 
a_{31} & a_{32} & a_{33} \\
\end{array}\right|=a_{11}*a_{22}*a_{33}+a_{12}*a_{23}*a_{31}+a_{13}*a_{21}*a_{32}-\\-a_{13}*a_{22}*a_{31}-a_{11}*a_{23}*a_{32}-a_{12}*a_{21}*a_{33}"

"x=\frac{D_x}{D},y=\frac{D_y}{D},z=\frac{D_Z}{D}"

"D=\det\left|
\begin{array}{ c c c }
5 & 6 & 7 \\ 
2 & 4 & 2\\ 
1 & 3 & 5 \\
\end{array}\right|=5*4*5+6*2*1+7*2*3-7*4*1-\\
-5*2*3-6*2*5=36"

"D_x=\det\left|
\begin{array}{ c c c }
40 & 6 & 7 \\ 
34 & 4 & 2\\ 
30 & 3 & 5 \\
\end{array}\right|=40*4*5+6*2*30+7*34*3-7*4*30-\\-40*2*3-6*34*5=-226"

"D_y=\det\left|
\begin{array}{ c c c }
5 & 40 & 7 \\ 
2 & 34 & 2\\ 
1 & 30 & 5 \\
\end{array}\right|=5*34*5+40*2*1+7*2*30-7*34*1-\\-5*2*30-40*2*5=412"

"D_z=\det\left|
\begin{array}{ c c c }
5 & 6 & 40 \\ 
2 & 4 & 34\\ 
1 & 3 & 30 \\
\end{array}\right|=5*4*30+6*34*1+40*2*3-40*4*1-\\-5*34*3-6*2*30=14"

"x=\frac{D_x}{D}=\frac{-226}{36}=-\frac{113}{18}=-6\frac{5}{18};y=\frac{D_y}{D}=\frac{412}{36}=\frac{103}{9}=11\frac{4}{9};z=\frac{D_z}{D}=\frac{14}{36}=\frac{7}{18}"

Answer :

"x=-6\frac{5}{18};y=11\frac{4}{9};z=\frac{7}{18}"